# Market Sensitivities and Growth Differentials Across Australian Housing Markets

### v2

@@ 2.5 The factor model and uncertainty in projections relative to the market
The model decomposes each regional index $\mu_{r}$ into a Market component (capturing long-run growth), a Mining component (capturing resource-sector cycles), and a Lifestyle component (capturing amenity-driven demand shifts), plus an idiosyncratic remainder. The Market loadings determine long-run growth differentials, while the remaining factors are thought of as bounded, representing socio-economic cycles.
@@
where $U$ is the national market index, $\delta_{PS}$ the Perth–Sydney spread, $\delta_{L}$ the Lifestyle spread, and $\epsilon_{r}$ an idiosyncratic disturbance. Although autocorrelated, we assume $\mathbb{E}[\delta_{PS}]=\mathbb{E}[\delta_{L}]=\mathbb{E}[\epsilon_{r}]=0$. We treat $(\beta_{r},\lambda_{r},\gamma_{r})$ as time-invariant constants, providing justification for this in Section 3.3.
@@
Let $\mathcal{H}_{t_{1}}$ be the information set at $t_{1}$ and define the horizon length as $h=t_{2}-t_{1}$. Since $X(t_{1})$ is known at $t_{1}$, the variance of the level change equals the $h$-step level forecast variance:
$\mathrm{Var}\!\big(X(t_{2})-X(t_{1})\,\big|\,\mathcal{H}_{t_{1}}\big)\;=\;\mathrm{Var}\!\big(X(t_{2})\,\big|\,\mathcal{H}_{t_{1}}\big).$
We therefore define the three horizon variances (“funnels”) as:
@@
We estimate this model using ARIMAX, with the Market, Mining and Lifestyle factors as external regressors and ARMA dynamics for the error term. The regional log real price indices $\mu_{r}(t)$ and the Market factor $U_{t}$ are nonstationary, whereas the Mining and Lifestyle spreads are approximately stationary, so we treat the factor regression in levels as describing a long-run relationship between $\mu_{r}(t)$ and $U_{t}$, with short-run dynamics captured by the ARMA error. In Section 3.4 we show that the estimated residuals $\varepsilon_{r,t}$ are bounded and display no visible trend, which is consistent with an approximately stationary error process.
Remark. The quantities $\sigma^{2}_{PS}(h)$, $\sigma^{2}_{L}(h)$, and $\sigma^{2}_{\epsilon}(h)$ are the $h$-step conditional level variances (forecast “funnels”) implied by the fitted dynamics of the spreads and the idiosyncratic component; loadings are fixed and are not projected.
@@ 3 Results
The three-factor model of price indexes of regions within Australia, with a market trend, absorbing the main trend and linking local growth to national price developments, alongside two mean-reverting trends, provides a framework for attributing regional price growth to the main factors, and compare regions in terms of their fundamental growth, captured in their loading on the Market factor. A central concern of this paper is the temporal stability of these factor loadings, particularly $\beta_{r}$, i.e. with respect to an expanding fit window. In addition, to assess the uncertainties and risks introduced by the remaining mean-reverting factors, we are interested in their forecast funnels over a future period: we estimate this by individual ARIMA models for each factor. Finally, We adopt an ARIMAX specification for the factor model of each individual city in Equation 3, allowing a good estimation of the regression coefficients and the construction of a forecast funnel for the autocorrelated remainder.
@@ 3.1 Factor model selection for cities
We estimate ARIMAX(p,0,q)(p,0,q) models in levels ($d=0$), with the unit-root absorbed by the Market factor $U_{t}$; Mining and Lifestyle enter as spreads, and the remainder $\epsilon_{r}(t)$ is treated as stationary. Although an ADF test on the raw Mining spread does not reject a unit root, this mainly reflects structural breaks around the mining boom; Appendix D shows that once break regimes are accounted for, the Mining factor is well described as a persistent but mean-reverting cycle, consistent with our levels ARIMAX specification.
Visual inspection shows no drift in $\epsilon_{r}(t)$ after accounting for the factors (Section 3.4, Figures 57). Non-seasonal orders $(p,q)\in\{0,1,2\}\times\{0,1,2\}$ are chosen by AICc (hurvich_tsai1989) with a seasonal MA(1) at $s=12$; low orders dominate (typically $(2,0,1)$ or $(2,0,0)$). Full selections and Ljung–Box diagnostics (ljung_box1978) are reported in Appendix Table 7.
LB12 (and LB24) denote Ljung–Box portmanteau $p$-values testing for no residual autocorrelation up to lag 12 (24), computed on the one-step-ahead prediction errors $e_{r}(t)=y_{r}(t)-\hat{y}_{r}(t\mid t-1)$ from each fitted ARIMAX (unstandardized). Degrees of freedom are not adjusted for estimated ARMA/seasonal parameters.
@@
We wish to include at least two factors, and now assess whether adding Lifestyle improves fit by comparing Market+Mining (2f) to Market+Mining+Lifestyle (3f). Inclusion is favored when $\Delta\mathrm{AICc}\leq-2$ (burnham_anderson2002). Ten of 13 regions meet this rule, with large gains for Melbourne (287.5-287.5), Rest of QLD (100.2-100.2), and Rest of NSW (98.3-98.3). Residual diagnostics are generally satisfactory (LB12), so we adopt the three-factor specification across the board for comparability and motivated by the PCA; see Table 4. For example, although the Rest of QLD does not pass the LB12 test under 3f (unlike 2f), we nonetheless see that $\Delta\mathrm{AICc}$ provides a large improvement. Residual variance also shrinks (not shown).
@@
Table 4: Two– vs three–factor ARIMAX in levels (Market+Mining vs Market+Mining+Lifestyle). Reported are AICc for each spec and $\Delta\mathrm{AICc}=\mathrm{AICc}_{\text{3f}}-\mathrm{AICc}_{\text{2f}}$ (negative favors three factors). Lifestyle is included when $\Delta\mathrm{AICc}\leq-2$; “Yes” indicates LB12 $\geq 0.05$.
@@ 3.2 The stability of the factor series
To estimate the uncertainties of Equation 4 in Section 2.5, we need to first obtain the forecast uncertainties of the individual Mining and Lifestyle factor time series (as our work is relative to the market, we do not need this for $U$): to this end, we employ an individual ARIMA model for each of the two factors. The ten-year ARIMA forecasts and 95% prediction intervals (“funnels”) for the factor spreads shown in Figure 2 demonstrate that both spread series, the Perth–Sydney “Mining” spread $\delta_{PS}$ and the “Lifestyle” spread $\delta_{L}$, are stationary processes well described by low–order ARIMA models in levels with a constant and no differencing required.
@@
Figure 2: Ten-year ARIMA forecasts and 95% prediction intervals (“funnels”) for the factor spreads: (a) Mining; (b) Lifestyle.
@@
Table 5 reports the implied 120–month–ahead standard deviations and their multiplicative 95% factors at the factor level. For Mining we obtain $\sigma_{10y}=0.27$, corresponding to a factor–level band of $x95=\exp(1.96\,\sigma)=1.70$; Lifestyle is tighter with $\sigma_{10y}=0.15$ and $x95=1.33$. In multiplicative terms, a 70% range for Mining and 33% range for Lifestyle at the ten-year horizon indicate moderate but non-negligible long-run uncertainty. Importantly, these values are factor-level uncertainties and must be scaled by regional loadings to obtain region-specific bands. In our regional scenario projections, the factor contributions enter quadratically via
@@
The stable funnels and whitened residuals validate our use of fixed loadings with uncertainty propagated through factor–specific forecast variances. While Mining exhibits larger factor-level uncertainty ($\sigma=0.27$ vs $\sigma=0.15$ for Lifestyle), whether it dominates a given region’s total uncertainty band depends critically on that region’s factor exposures $(\lambda_{r},\gamma_{r})$. For instance, Perth’s large Mining loading amplifies Mining uncertainty substantially. This decomposition underpins the segmented uncertainty bars and the $x95_{\text{total}}$ values in the next section (Figure 8 and Table 6).
@@ 3.3 The factor loadings and their stability
To test the assumption of our scenario framework that the factor loadings $(\beta_{r},\lambda_{r},\gamma_{r})$ remain stable over time, we estimate the model of Section 2.5 on a sequence of expanding windows, each starting in January 1995 and ending at a point between January 2008 and December 2024, advancing in three-month steps. We hold the ARIMAX orders fixed at the values selected in Section 3.1 (Table 7): these are specifications in levels with seasonal MA(1) at $s=12$. Figures 34 plot the resulting coefficient paths $(\beta_{r},\lambda_{r},\gamma_{r})$ against the window endpoint.
The loadings exhibit remarkable stability over time, albeit with some temporary disturbances or spikiness. For the major cities, the market sensitivity $\beta_{r}$ typically varies within a range of $\pm 0.05$$0.10$ around its median, with no discernible trend [^7] For Brisbane, we interpret the downward excursion for the most recent loadings as a temporary glitch., and the spread loadings $(\lambda_{r},\gamma_{r})$ show similarly bounded variation consistent with our interpretation of the spreads as stationary processes.
Brief excursions do occur, most notably for Rest of NSW around 2021–2022, where $\beta_{r}$ spikes temporarily during the post-pandemic adjustment, but these disturbances are transient and revert within a few quarters. Importantly, there is no evidence of trending behavior in any of the loadings, which would signal decreased utility of the factor structure for future expectations.
To guard against the influence of transient noise in any single window, we calculate each coefficient by its median across window endpoints from 2014 to 2024 (we select 2014, as a shorter sample may induce estimation instability without yielding further insights). These median loadings are reported in Table 6. Melbourne stands out with the highest Market beta ($\beta_{r}>1$), Brisbane has an elevated value, while Sydney and the ACT are close to unity and regional areas show dampening ($\beta_{r}<1$) relative to the national market. The Mining and Lifestyle loadings similarly align with qualitative regional narratives (e.g., Perth’s strong Mining exposure, Melbourne’s negative Lifestyle loading). As a robustness check, Appendix Table 10 reports full-sample ARIMAX estimates and 95% confidence intervals for the factor loadings; these confirm that Melbourne’s Market beta is clearly above one, whereas the Market betas for Sydney, Brisbane and the ACT are relatively close to unity.
@@ 3.4 Examples of price indexes, their decompositions, and relative valuation
Figures 57 illustrate the factor decomposition for three representative markets: the Rest of NSW, ARIMAX orders $(2,0,0)$, and the major city areas Greater Sydney and Greater Perth, each ARIMAX orders $(2,0,1)$. Each is estimated with seasonal orders $(0,0,1)_{12}$. The top panel of each figure overlays the observed monthly log house price index with cumulative factor approximations added sequentially (Market; Market+Mining; Market+Mining+Lifestyle) and the ARIMAX fitted mean. As factors are added, the approximation tightens; with all three factors the fitted mean is nearly indistinguishable from the observed index at the plotting scale.
The bottom panel reports the remainder, defined as the ARIMAX fitted mean minus the three-factor approximation. Across all three markets this remainder exhibits no visual drift, consistent with the factors capturing the low-frequency co-movement, and shows a gentle undulating pattern consistent with an ARMA component plus seasonal MA(1) at 12 months, supporting the $(2,0,1)$ with $(0,0,1)_{12}$ specification. In addition, the Ljung–Box $p$-values in Table 4 are generally above 0.05 for the selected specifications, indicating no evidence of remaining autocorrelation up to 12 lags.
@@
The Rest of NSW (Figure 5) is commonly characterized as a lifestyle market with limited direct exposure to mining. This is reflected in its decomposition: despite the presence of mining regions within NSW (e.g. Singleton), the Mining factor makes little difference to the Market-only approximation (in fact, the loading is slightly negative, indicating a Sydney influence), whereas adding the Lifestyle factor explains a large share of the overall growth pattern. As in other lifestyle regions, this imparts a more cyclical profile: a sharp run-up into the mid-2000s peak (circa 2004), a prolonged plateau through roughly 2012, and renewed growth that accelerates during the COVID-19 period beginning in 2020.
@@
The Sydney market (Figure 6) is not strongly influenced by Lifestyle; instead, adding the Mining factor materially improves the fit, reflecting Sydney’s negative Mining loading ($\lambda_{\text{Sydney}}=-0.43$). This pattern is already apparent in sijp_francke_ree2024, and remains somewhat unexplained. One context is Australia’s terms of trade (ToT). Using national quarterly data, tumbarello_wang2010 find that positive ToT shocks raise Australian real house prices in the long run, while downes_et_al_rba_report2014 provide a detailed analysis of the mining boom.
In our decomposition, this ToT/mining-boom impulse is absorbed into the Market factor, and the Mining spread models local variation around this common influence: Figure 6 shows Sydney underperforming its Market-factor baseline through the mid-2000s boom phase, consistent with $\lambda_{\text{Sydney}}<0$. Sydney growth during 2004–2006 is quite flat. This could reflect a reduced sensitivity of Sydney to the mining boom, or a net negative effect of the resource cycle on Sydney price growth via channels such as internal migration and local demand and supply conditions. Australian evidence suggests that house prices have become increasingly sensitive to mortgage rates, with the largest marginal effects in Sydney and Melbourne (otto2007). In addition, Sydney and Melbourne are more exposed than the other capitals to foreign real-estate investment, and comparative work on large cities emphasises the growing role of global investors and financial conditions in driving price dynamics in major urban housing markets. [^8] See, for example, the discussion of foreign investors and global financial conditions in large cities in Nijskens2019HotProperty. Insofar as the mining boom pushed the Australian dollar to historically high levels and was associated with higher interest rates, Sydney housing would have become more expensive to foreign-currency buyers precisely when domestic financing conditions were tightening. This combination could plausibly have dampened Sydney’s response to the mining boom, with a similar but weaker effect in Melbourne.
@@
A pattern converse to Sydney appears for Perth (Figure 7). The mining boom lifts prices relative to Perth’s Market factor (consistent with a strong positive loading, $\lambda_{\text{Perth}}=0.60$), with a subsequent decline after the boom’s end in 2013 and a reversal beginning around 2020 marked by a sharp increase thereafter. Unlike the Lifestyle factor, the Mining factor explains a substantial share of the deviation of Perth’s index from the Market path, reflecting the mining sector’s influence on employment, population flows, and the local economy.

### v3

@@ 2.5 The factor model and uncertainty in projections relative to the market
The model decomposes each regional index $\mu_{r}$ into a Market component (capturing long-run growth), a Mining component (capturing resource-sector cycles), and a Lifestyle component (capturing amenity-driven demand shifts), plus a city-specific idiosyncratic remainder. The Market loadings determine long-run growth differentials, while the remaining factors are thought of as bounded, representing socio-economic cycles.
@@
where $U$ is the national market index, $\delta_{PS}$ the Perth–Sydney spread, $\delta_{L}$ the Lifestyle spread, and $\epsilon_{r}$ an idiosyncratic disturbance. Although autocorrelated, we assume $\mathbb{E}[\delta_{PS}]=\mathbb{E}[\delta_{L}]=\mathbb{E}[\epsilon_{r}]=0$. We treat $(\beta_{r},\lambda_{r},\gamma_{r})$ as time-invariant constants, providing justification for this in Section 3.3. The coefficients $(\beta_{r},\lambda_{r},\gamma_{r})$ represent regional sensitivities to the three factors, analogous to loadings on the corresponding principle components.
@@
Let $\mathcal{H}_{t_{1}}$ be the information set at $t_{1}$ and define the horizon length as $h=t_{2}-t_{1}$. Since $X(t_{1})$ is known at $t_{1}$ for a series $X$, the variance of the level change equals the $h$-step level forecast variance: $\mathrm{Var}\!\big(X(t_{2})-X(t_{1})\,\big|\,\mathcal{H}_{t_{1}}\big)\;=\;\mathrm{Var}\!\big(X(t_{2})\,\big|\,\mathcal{H}_{t_{1}}\big).$ We therefore define the three horizon variances (“fans”) as:
@@
Unlike many regressions of returns in financial time series, the city-specific idiosyncratic disturbance $\varepsilon_{r,t}$ is autocorrelated (this would even be the case if we were to model housing index returns). Therefore, we do not estimate Eq. 3 with OLS. Instead, we model the disturbance as a low-order ARMA process,
$\varepsilon_{r,t}=\sum_{i=1}^{p_{r}}\phi_{r,i}\,\varepsilon_{r,t-i}+u_{r,t}+\sum_{j=1}^{q_{r}}\theta_{r,j}\,u_{r,t-j},\qquad u_{r,t}\sim\text{i.i.d. }(0,\sigma_{r}^{2}),$ (5)
[…]
The regional log real price indices $\mu_{r}(t)$ and the Market factor $U_{t}$ are nonstationary, whereas the Mining and Lifestyle spreads are approximately stationary (Appendix D), so we treat the factor regression in levels as describing a long-run relationship between $\mu_{r}(t)$ and $U_{t}$, with the bounded dynamics captured by the remaining factors an $\epsilon$. In Section 3.4 we show that the $\epsilon$ are bounded and display no visible trend, which is consistent with an approximately stationary process.
Remark. The quantities $\sigma^{2}_{PS}(h)$, $\sigma^{2}_{L}(h)$, and $\sigma^{2}_{\epsilon}(h)$ are the $h$-step conditional level variances (forecast “fans”) implied by the fitted dynamics of the spreads and the idiosyncratic component; loadings are fixed and are not projected.
@@ 3 Results
City price indexes are modeled with three-factors and a city-specific component ($\epsilon$). A concern of this paper is the temporal stability of the factor loadings, particularly $\beta_{r}$, i.e. with respect to an expanding fit window. In addition, to assess the uncertainties and risks introduced by the remaining mean-reverting factors, we are interested in their forecast fans over a future period: we estimate this by individual ARIMA models for each factor. Finally, We adopt an ARIMAX specification for the factor model of each individual city in Equation 3, allowing an estimation of the regression coefficients and the construction of a forecast funnel for the autocorrelated remainder.
@@ 3.1 Factor model selection for cities
We estimate ARIMAX(p,0,q)(p,0,q) models for Equation 3 in levels ($d=0$), with the unit-root absorbed by the Market factor $U_{t}$; Mining and Lifestyle enter as spreads, and the remainder $\epsilon_{r}(t)$ is treated as stationary. Although an ADF test on the raw Mining spread does not reject a unit-root, this mainly reflects structural breaks around the mining boom; Appendix D shows that once break regimes are accounted for, the Mining factor is well described as a persistent but mean-reverting cycle, consistent with our levels ARIMAX specification.
We will see that visual inspection shows no drift in $\epsilon_{r}(t)$ after accounting for the factors (Section 3.4, Figures 57). Non-seasonal orders $(p,q)\in\{0,1,2\}\times\{0,1,2\}$ are chosen by AICc (hurvich_tsai1989) with a seasonal MA(1) at $s=12$; low orders dominate (typically $(2,0,1)$ or $(2,0,0)$). Full selections and Ljung–Box diagnostics (ljung_box1978) are reported in Appendix Table 7. LB12 (and LB24) denote Ljung–Box portmanteau $p$-values testing for no residual autocorrelation up to lag 12 (24), computed on the one-step-ahead prediction errors $e_{r}(t)=y_{r}(t)-\hat{y}_{r}(t\mid t-1)$ from each fitted ARIMAX (unstandardized). Degrees of freedom are not adjusted for estimated ARMA/seasonal parameters.
@@
We wish to include at least two factors, and now assess whether adding Lifestyle improves fit by comparing Market+Mining (2f) to Market+Mining+Lifestyle (3f). Inclusion is favored when $\Delta\mathrm{AICc}\leq-2$ (burnham_anderson2002). Ten of 13 regions meet this rule, with large gains for Melbourne (287.5-287.5), Rest of QLD (100.2-100.2), and Rest of NSW (98.3-98.3). Residual diagnostics on $e_{r,t}$ are generally satisfactory (LB12) with some exceptions, so we adopt the three-factor specification across the board for comparability and motivated by the PCA; see Table 4. For example, although the Rest of QLD does not pass the LB12 test under 3f (unlike 2f), we nonetheless see that $\Delta\mathrm{AICc}$ provides a large improvement. Residual variance also shrinks (not shown).
@@
Table 4: Two– vs three–factor ARIMAX in levels (Market+Mining vs Market+Mining+Lifestyle). Reported are AICc for each spec and $\Delta\mathrm{AICc}=\mathrm{AICc}_{\text{3f}}-\mathrm{AICc}_{\text{2f}}$ (negative favors three factors). Lifestyle is included when $\Delta\mathrm{AICc}\leq-2$; The asterisk in “Yes” indicates LB12 $\geq 0.05$.
@@ 3.2 The stability of the factor series
To estimate the uncertainties of Equation 4 in Section 2.5, we need to first obtain the forecast uncertainties of the individual Mining and Lifestyle factor time series (as our work is relative to the market, we do not need this for $U$): to this end, we employ an individual ARIMA model for each of the two factors [^6] This is a separate model for each of the 2 factors, not to be confused with the city-specific ARIMAX estimations of the factor model..
@@
Figure 2: Ten-year ARIMA forecasts and 95% prediction intervals (“fans”) for the factor spreads: (a) Mining; (b) Lifestyle.
@@
The ten-year ARIMA forecasts and 95% prediction intervals for the factor spreads shown in Figure 2 demonstrate that both spread series remain reasonably bounded in the future. We use these ARIMA-based fans as inputs to the uncertainty estimates in Equation 4, interpreting them as conditional on the maintained assumption that Mining and Lifestyle are stationary (or highly persistent) spreads.
@@
Table 5 reports the implied 120–month–ahead standard deviations and their multiplicative 95% factors at the factor level. For Mining we obtain $\sigma_{10y}=0.27$, corresponding to a factor–level band of $x95=\exp(1.96\,\sigma)=1.70$; Lifestyle is tighter with $\sigma_{10y}=0.15$ and $x95=1.33$. In multiplicative terms, a 70% range for Mining and 33% range for Lifestyle at the ten-year horizon indicate moderate long-run uncertainty that we can use in our city-specific estimations. Importantly, these values are factor-level uncertainties and must be scaled by regional loadings to obtain region-specific bands. In our regional scenario projections, the factor contributions enter quadratically via
@@
The stable fans and whitened residuals validate our use of fixed loadings with uncertainty propagated through factor–specific forecast variances.
@@ 3.3 The factor loadings and their stability
To test the assumption of our scenario framework that the factor loadings $(\beta_{r},\lambda_{r},\gamma_{r})$ remain stable over time, we estimate the model of Section 2.5 on a sequence of expanding windows, each starting in January 1995 and ending at a point we select for testing. We select endpoints between January 2008 and December 2024, advancing in three-month steps. We hold the ARIMAX orders fixed at the values selected in Section 3.1 (Table 7): these are specifications in levels with seasonal MA(1) at $s=12$. Figures 34 plot the resulting coefficient paths $(\beta_{r},\lambda_{r},\gamma_{r})$ against the window endpoint.
The loadings exhibit stability over time, albeit with some temporary disturbances or spikiness. For the major cities, the market sensitivity $\beta_{r}$ typically varies within a range of $\pm 0.05$$0.10$ around its median, and the spread loadings $(\lambda_{r},\gamma_{r})$ show similarly bounded variation consistent with our interpretation of the spreads as stationary processes. Importantly, there is no discernible trend in any of the loadings [^8] For Brisbane, we interpret the downward excursion for the most recent loadings as a temporary glitch., which would signal decreased utility of the factor structure for future expectations.
To guard against the influence of transient noise in any single window in our conclusions, we calculate each coefficient by its median across window endpoints from 2014 to 2024 (we select 2014, as a shorter sample may induce estimation instability without yielding further insights). These median loadings are reported in Table 6. Melbourne stands out with the highest Market beta ($\beta_{r}>1$), Brisbane has an elevated value, while Sydney and the ACT are close to one and regional areas show dampening ($\beta_{r}<1$) relative to the national market. The Mining and Lifestyle loadings similarly align with qualitative regional narratives (e.g., Perth’s strong Mining exposure, Melbourne’s negative Lifestyle loading). As a robustness check, Appendix Table 10 reports full-sample ARIMAX estimates and 95% confidence intervals for the factor loadings; these confirm that Melbourne’s Market beta is clearly above one in the statistical sense, whereas the Market betas for Sydney, Brisbane and the ACT are relatively close to one.
@@ 3.4 Examples of price indexes, their decompositions, and relative valuation
It is instructive to examine the price index behavior of a few individual price indexes. Figures 57 illustrate the factor decomposition for three representative markets: the Rest of NSW, with ARIMAX orders $(2,0,0)$, and the major city areas Greater Sydney and Greater Perth, each with ARIMAX orders $(2,0,1)$. Each is estimated with seasonal orders $(0,0,1)_{12}$. The top panel of each figure overlays the observed monthly log house price index with cumulative factor approximations added sequentially (Market; Market+Mining; Market+Mining+Lifestyle) and the ARIMAX fitted mean. As factors are added, the approximation tightens; with all three factors the fitted mean is nearly indistinguishable from the observed index at the plotting scale.
The bottom panel reports the remainder $\epsilon_{r,t}$, defined as the ARIMAX fitted mean minus the three-factor approximation. Across all three markets this remainder exhibits no visual drift, consistent with the factors capturing the low-frequency co-movement, and shows a gentle undulating pattern consistent with an ARMA component plus seasonal MA(1) at 12 months, supporting the $(2,0,1)$ with $(0,0,1)_{12}$ specification. In addition, the Ljung–Box $p$-values in Table 4 are generally above 0.05 for the selected specifications, indicating no evidence of remaining autocorrelation up to 12 lags.
@@
The Rest of NSW (Figure 5) is commonly characterized as a lifestyle market with limited direct exposure to mining. This is reflected in the decomposition of its log price index: despite the presence of mining regions within NSW (e.g. Singleton), the Mining factor makes little difference to the Market-only approximation (in fact, the loading is slightly negative, indicating a Sydney influence), whereas adding the Lifestyle factor (green series) explains a large share of the overall growth pattern. As in other lifestyle regions, this imparts a more cyclical profile: a sharp run-up into the mid-2000s peak (circa 2004), a prolonged plateau through roughly 2012, and renewed growth that accelerates during the COVID-19 period beginning in 2020.
@@
The Sydney log price index (Figure 6) is not strongly influenced by Lifestyle; instead, adding the Mining factor (orange series) materially improves the fit compared to the market-only approximation (blue), reflecting Sydney’s negative Mining loading ($\lambda_{\text{Sydney}}=-0.43$). This pattern is apparent in sijp_francke_ree2024, and remains somewhat unexplained. One context is Australia’s terms of trade (ToT). Using national quarterly data, tumbarello_wang2010 find that positive ToT shocks raise Australian real house prices in the long run, while downes_et_al_rba_report2014 provide a detailed analysis of the mining boom.
In our decomposition, the Australia-wide effect of this ToT/mining-boom impulse is absorbed into the Market factor, and the Mining spread models local variation around this common influence: Figure 6 shows Sydney underperforming its Market-factor baseline through the mid-2000s boom phase, consistent with $\lambda_{\text{Sydney}}<0$. This growth during 2004–2006 is quite flat, and could reflect a reduced sensitivity of Sydney to the mining boom, or a net negative effect of the resource cycle on Sydney price growth via channels such as internal migration and local demand and supply conditions. Australian evidence suggests that house prices have become increasingly sensitive to mortgage rates, with the largest marginal effects in Sydney and Melbourne (otto2007). In addition, Sydney and Melbourne are more exposed than the other capitals to foreign real-estate investment, and comparative work on large cities emphasises the growing role of global investors and financial conditions in driving price dynamics in major urban housing markets. [^9] See, for example, the discussion of foreign investors and global financial conditions in large cities in Nijskens2019HotProperty. Insofar as the mining boom pushed the Australian dollar to historically high levels and was associated with higher interest rates, Sydney housing would have become more expensive to foreign-currency buyers precisely when domestic financing conditions were tightening. This combination could plausibly have dampened Sydney’s response to the mining boom, with a similar but weaker effect in Melbourne.
@@
A pattern converse to Sydney appears for Perth (Figure 7). The mining boom lifts prices relative to Perth’s Market factor (consistent with a strong positive loading, $\lambda_{\text{Perth}}=0.60$), with a subsequent decline after the boom’s end in 2013 and a reversal beginning around 2020 marked by a sharp increase thereafter. Unlike the Lifestyle factor, the Mining factor explains a substantial share of the deviation of Perth’s index from the Market path, reflecting the mining sector’s influence on employment, population flows, incomes, and the local economy.